A boy of mass 25.0 kg skateboards down a quarter pipe of radius 3.00 m. He starts from rest and there is no friction. a)Find his speed at the bottom b)Find the normal force at the bottom Example 7.5: if the speed at the bottom is instead 6.00 m/s, what was the work done by friction?
Answer #1

a)

using energy conservation

0.5 m v^2 = m g h

0.5 v^2 = 9.8* 3

v = 7.668 m/s

======

b)

W = change in kE

W = 0.5* 25 * ( 6^2 - 7.668^2)

W = - 285 J